Solve the system of equations using the addition (elimination) method.
If the answer is a unique solution, present it as an ordered pair: (x, y). If not, specify whether the answer is "no solution" or "infinitely many solutions" and state how you arrived at that conclusion.
4x + 10y = 2
3x + 5y = 5

Answer by stanbon(75887)

4x + 10y = 2
3x + 5y = 5
--------------
Multiply 2nd by -2 to get:
4x + 10y = 2
-6x -10y = -10
-------------------
Add to get:
-2x = -8
x = 4
-------------------
Substitute to solve for "y":
4*4 + 10y = 2
10y = -14
y = -7/5
=================
Cheers,
Stan H.